OpenStudy (anonymous):

4. Multiply. give your answer in standard form. (2n^2 - 7n + 2)(3n + 1) (1 point) 6n^3 - 23n62 - n +2 6n^3 - 19n^2 - n + 2 6n^3 - 19n^2 - 13n + 2 6n^3 + 23n^2 + 13n + 2

jagr2713 (jagr2713):

what do you think it is

OpenStudy (anonymous):

i have no idea @jagr2713

jagr2713 (jagr2713):

7-4=3 3+3=6

jagr2713 (jagr2713):

so 6n^3

jagr2713 (jagr2713):

7*3=21 21-2=19

jagr2713 (jagr2713):

6n^3-19n^2

jagr2713 (jagr2713):

so what can you eliminate

OpenStudy (anonymous):

the frist one n last one

jagr2713 (jagr2713):

ok so its leaves b and c so

OpenStudy (anonymous):

so what is it

jagr2713 (jagr2713):

what is 1

OpenStudy (anonymous):

1

jagr2713 (jagr2713):

x=1 so whats the answer

OpenStudy (anonymous):

b?

jagr2713 (jagr2713):

Yes nice

OpenStudy (anonymous):

ok kool thanx

jagr2713 (jagr2713):

any other questions like that? and you're welcome

OpenStudy (anonymous):

the area of a rectangle painting is given by the trrinomial x^2+2x-15. What are the possible dimentions? use factoring.

jagr2713 (jagr2713):

answer choice??

OpenStudy (anonymous):

(x+5) and (x-3) (x+5) and (x+3) (x-5) and (x-3) (x+5)and (x+3)

jagr2713 (jagr2713):

x^2+2x=3x 15/3=5 so its (x+5) right

jagr2713 (jagr2713):

so what can you eliminate

OpenStudy (anonymous):

the last two the last one is supposed to be (x-5)

jagr2713 (jagr2713):

what?

OpenStudy (anonymous):

its either the first one or the second

jagr2713 (jagr2713):

x^2 +2=3 b/c it has - what would the answer be

OpenStudy (anonymous):

b?

jagr2713 (jagr2713):

Nice again

OpenStudy (anonymous):

kool

jagr2713 (jagr2713):

wait no no its not

jagr2713 (jagr2713):

hint it has -

jagr2713 (jagr2713):

so the answer is A

OpenStudy (anonymous):

ok thanx

jagr2713 (jagr2713):

and where are they unanswered questions

OpenStudy (anonymous):

alot bc its a 41 question thing that i have no idea how to do so yea...

jagr2713 (jagr2713):

ok i can help with a few more

OpenStudy (anonymous):

ok (6x^3-1x^2+1) / (2x +1)

jagr2713 (jagr2713):

answer choice

jagr2713 (jagr2713):

6x*2=12 x+x^3=4 12x^4

jagr2713 (jagr2713):

thats the first one

OpenStudy (anonymous):

-3x^2+2x-1 3x^2 -2x+1 -3x^2+2x-1 3x^3+2x-1

jagr2713 (jagr2713):

oh ok wait

jagr2713 (jagr2713):

can you send pic of the question please

OpenStudy (anonymous):

not really

jagr2713 (jagr2713):

its this (6x^3-1x^2+1) / (2x +1)

OpenStudy (anonymous):

yep

jagr2713 (jagr2713):

ohoho its divide ok lets do this

jagr2713 (jagr2713):

6-2-1=3x^2 right so what can you eliminate

OpenStudy (anonymous):

a n c

jagr2713 (jagr2713):

ok 6-2-2-1=2x so whats the answer

jagr2713 (jagr2713):

hint: positive 2x

OpenStudy (anonymous):

d?

jagr2713 (jagr2713):

opps sorry i am messing up i mean negitive 2x so its B

OpenStudy (anonymous):

ok

jagr2713 (jagr2713):

:D i love doing these

OpenStudy (anonymous):

ewwww i hate algebra

jagr2713 (jagr2713):

any more questions

OpenStudy (anonymous):

y= x^2 +5x -9 y=2x+1 (2,-9) n (-5,5) (2,5) n (-5,-5) (2,5) n (-5,-9) no solution

jagr2713 (jagr2713):

i hate these ones but ok

jagr2713 (jagr2713):

Its c

OpenStudy (anonymous):

ok nice thanks

jagr2713 (jagr2713):

what are you on flvs

OpenStudy (anonymous):

ummm no whats that

jagr2713 (jagr2713):

oh forget it and any more

OpenStudy (anonymous):

no

jagr2713 (jagr2713):

Ok it was nice answering your questions

OpenStudy (anonymous):

thanx for answering them

jagr2713 (jagr2713):

you're welcome